# Driving with Prior Maps: Unified Vector Prior Encoding for Autonomous Vehicle Mapping

### v3

@@ Abstract
High-Definition Maps (HD maps) are essential for the precise navigation and decision-making of autonomous vehicles, yet their creation and upkeep present significant cost and timeliness challenges. The online construction of HD maps using on-board sensors has emerged as a promising solution; however, these methods can be impeded by incomplete data due to occlusions and inclement weather. This paper proposes the PriorDrive framework to addresses these limitations by harnessing the power of prior maps, significantly enhancing the robustness and accuracy of online HD map construction. Our approach integrates a variety of prior maps, such as OpenStreetMap’s Standard Definition Maps (SD maps), outdated HD maps from vendors, and locally constructed maps from historical vehicle data. To effectively encode this prior information into online mapping models, we introduce a Hybrid Prior Representation (HPQuery) that standardizes the representation of diverse map elements. At the core of PriorDrive is the Unified Vector Encoder (UVE), which employs hybrid prior embedding and a dual encoding mechanism to process vector data. Furthermore, we propose a segment-level and point-level pre-training strategy that enables the UVE to learn the prior distribution of vector data, thereby improving the encoder’s generalizability and performance. Through extensive testing on the nuScenes, Argoverse 2 and OpenLane-V2, we demonstrate that PriorDrive is highly compatible with various online mapping models and substantially improves map prediction capabilities. The integration of prior maps through the PriorDrive framework offers a robust solution to the challenges of single-perception data, paving the way for more reliable autonomous vehicle navigation.
@@
Figure 1: Overview of PriorDrive. Our unified vector encoder (UVE) directly encodes diverse vector prior maps and seamlessly integrates them into existing online mapping frameworks. This integration enhances the final predictions, making them more complete and accurate than those generated without prior information.
@@
footnotetext: Work done during the internship at Amap, Alibaba Group.footnotetext: Equal contribution.footnotetext: Corresponding author.
1 Introduction
@@
High-Definition Maps (HD maps) are essential for the accurate navigation and decision-making of autonomous vehicles, providing detailed vectorized representations of road elements such as lane dividers, pedestrian crossings, and road boundaries [6, 8]. Despite their significance, the traditional methods for creating and maintaining HD maps are often costly and labor-intensive. These methods can result in outdated maps that struggle to keep pace with the rapidly changing urban environments. In response, there is increasing interest in online HD map construction [27, 5, 38], where maps are generated in real-time using on-board sensors, such as panoramic RGB cameras and LiDAR point clouds. Although this approach reduces costs, it also introduces new challenges, particularly concerning incomplete and error-prone data caused by environmental occlusions or inclement weather conditions.
In this paper, we introduce PriorDrive, a novel framework designed to overcome these challenges by leveraging prior maps. Specifically, PriorDrive incorporates three types of prior maps: Standard Definition Maps (SD maps), existing but outdated HD maps, and online historical prediction maps, as illustrated in Figure 1. SD maps provide essential long-distance centerline skeletons, which are crucial for maintaining road-level connectivity. Existing HD maps (HD map-EX), despite their high accuracy, often fail to reflect current road conditions due to infrequent updates. Nevertheless, these maps offer valuable prior information for HD map construction. Additionally, historical prediction maps, which are often overlooked, can provide insights into current road conditions by reflecting the latest observations from previous perceptions. As vehicles traverse the same locations multiple times under varying conditions, the accumulated information can be used to enhance the quality of the current local map through iterative refinement.
Recent studies [35, 13, 29] have attempted to use prior maps to supplement perception data, but they face several limitations. For instance, NMP [35] relies on a neural prior that cannot take advantage of existing offline maps and can only be updated through model training, leading to poor transferability to new environments. P-MapNet [13] incorporates vector maps into the pipeline by converting them into rasterized representations, which are lossy and redundant, lacking the detail necessary to capture vectorized instance-level information and express the type and direction of map elements effectively. MapEX [29], on the other hand, focuses on existing HD maps but simulates outdated maps by introducing artificial offsets and erasing map elements, an approach that risks leaking ground truth data and fails to accurately represent the differences between simulated and actual outdated HD maps. These actual outdated maps, due to their low update frequency, struggle to reflect real-time changes in road structures, making them unreliable for current navigation needs.
In addition to constructing accurate online HD maps, the autonomous driving scenario requires understanding the topology in road scenes (i.e., lane connectivity and correspondence with traffic elements), which is called topology reasoning task. Since connected SD maps include crucial information for road topology, we also apply PriorDrive to this task to enhance the performance of topology reasoning.
Given that various maps are typically stored as vectors due to their low redundancy and universal applicability, it’s crucial to develop a robust vector encoding model. However, designing such a model is challenging due to the variable length, sparsity, and discreteness of vector data. To address this and enhance current perception results by integrating diverse vector prior maps into various mapping models, we propose a unified vector encoder (UVE). The UVE can encode a wide range of vector data types, and we introduce pre-training paradigms that allow model to fully learn the prior distribution of vector maps. Our experiments on nuScenes, Argoverse 2 and OpenLane-V2 demonstrate that our PriorDrive is plug-and-play and can be seamlessly applied to various online mapping models and vector tasks, resulting in significant performance improvements.
@@
  • We introduce a unified vector encoder that effectively captures diverse vector attributes and enables encoding of variable-length vectors into fixed-length features through Hybrid prior embedding and dual encoding mechanism.
@@
  • We first proposed a pre-training paradigm for vector data, which can effectively learn the prior distribution of vector data and help reduce noise in historical maps through point-level and segment-level pre-training.
@@
  • We propose a hybrid prior representation (HPQuery) to represent all elements and a PriorDrive framework with vector prior maps to addresses the limitations of single-perception. Our comprehensive evaluation demonstrates that PriorDrive is plug-and-play and significantly enhances the performance of online mapping models.
2 Related work
2.1 Online vectorized HD map construction
@@
Online vectorized HD map construction was initially treated as a segmentation task [15, 18]. To construct vectorized HD map, HDMapNet [15] requires complex post-processing of pixel-level rasterized maps. VectorMapNet [25] introduced a coarse-to-fine, two-stage network leveraging keypoint representations. MapTR [19] advanced this concept by modeling point sets in a permutation-equivalent manner using a DETR-like [2] one-stage network. Subsequent studies [41, 39, 36, 9, 12, 40, 3] have introduced various insightful approaches for further improvement. However, these methods rely solely on single-source perception data from vehicle-mounted sensors, limiting their effectiveness in challenging scenarios such as occlusion or inclement weather conditions.
@@
The release of OpenLane-V2 [31] dataset has sparked growing interest in topology reasoning within driving scenes. This task focuses on recognizing the topologies among lanes and lanes with traffic elements. TopoNet [17] leverages GNN to update lane representation and lane topology. TopoMLP [34] utilizes PETR [24] for centerline detection and employs simple MLPs to predict the relationships. In this paper, we also apply PriorDrive to topological inference tasks, introducing interconnected SD Map as an additional input to support topology reasoning.
2.2 Online mapping based on prior maps
@@
Recent research has explored the integration of prior maps to enhance the performance of online mapping models. P-MapNet [13] encodes Standard Definition Maps (SD map) as an additional conditional branch and employs a masked autoencoder to capture the prior distribution of HD maps. NMP [35] introduces a global neural map prior that self-updates, thereby improving the performance of local map inference. MapEX [29] categorizes existing maps into three types and refines the query-based map estimation model’s matching algorithm. Despite these advances, many studies overlook the potential of online historical prediction maps as a source of prior information. Our work addresses this gap by leveraging various existing prior maps in their native vector form to enhance current perception data.
2.3 Pretrained methods based on mask modeling
@@
In the fields of NLP and CV, masked pre-training has proven to be an effective strategy for self-supervised representation learning. In NLP, models like BERT [4, 23, 14] use masked language modeling to predict randomly masked tokens within a bidirectional text context. Similarly, methods such as MAE [11, 16, 21] mask random patches of input images and reconstruct them based on the remaining unmasked patches. Diverging from these approaches, which are predominantly tailored for text and image data, we introduce a pre-training paradigm specifically designed for vector data. Our method aims to learn the prior distribution of vector representations, offering a new avenue for pre-training in the context of vectorized data.
3 Proposed method: PriorDrive
Figure 2: Our Integration Method. For query-based models, we enable interaction between the prior features and queries at both instance-level and point-level, ensuring a more nuanced integration. Hierarchical query comes from MapTRv2 [20]. For the other models, we integrate prior features directly with the BEV features.
3.1 Problem formulation
@@
The goal of online mapping is to construct a local HD map using on-board sensor observations, such as the surrounding camera images. The online mapping process typically involves two main modules: the Bird’s-Eye View (BEV) encoder and the map decoder. The BEV encoder generally includes a Perspective View (PV) feature extractor and a module for transforming PV features into BEV features. The map decoder usually comprises a Fully Convolutional Network [28] or a query-based module similar to DETR [2], which outputs either rasterized perception results or vectorized map elements.
@@
where $\mathcal{P}$ represents the predicted map elements, $Q$ denotes the learnable query, and $opt$ indicates that this learnable query is optional, which is required in query-based models but not in other models.
@@
Our proposed UVE serves as a unified vector encoder that can directly encode various vector data information, such as the position ($x$, $y$), direction ($v_{x}$, $v_{y}$), and type of points ($c$). Using $p=\left[x,y,v_{x},v_{y},c\right]$ to represent the point, a vector $v$ typically consists of multiple ordered sets of points, $v=\left[p_{1},p_{2},\ldots,p_{n}\right]$, where $n$ varies. The map prior $M_{prior}$ consists of multiple vectors $M_{prior}=\{v_{1},v_{2},\ldots,v_{m}\}$, with $m$ being variable. Using $E_{uve}$ to represent the UVE model, the extraction process of the features of the prior map, denoted as $f_{prior}$, can be represented as:
@@
Online mapping based on prior maps.
We proposed four methods in total to integrate prior map features into online mapping models. For the methods without a learnable query $Q$, such as HDMapNet [15], we directly reshape the prior feature $f_{prior}$ and use $Deconv$ upsampling to match the shape of the BEV feature $f_{bev}$. Then connect it with the BEV feature $f_{bev}$ and aligned by $Conv$, as shown in Figure 2 left.
@@
For methods based on learnable query $Q$, such as the vectorized model MapTR series [20], we propose HPQuery of three operations: addition, replacement, and concatenation, as shown is Figure 2 right. Specifically, $Q$ is typically composed of two parts, $Q=\{q_{ij}\}_{i=0,j=0}^{m,n}=\{q_{i}^{ins}+q_{j}^{pt}\}_{i=0,j=0}^{m,n}$, $q_{i}^{ins}$ and $q_{i}^{pt}$ representing the learnable queries at the instance-level and point-level respectively. The features $f_{prior}$ of prior maps are also composed of instance-level features and point-level features, $f_{prior}=\{(f_{i}^{ins},f_{j}^{pt})\}_{i=0,j=0}^{m^{\prime},n^{\prime}}$. So we interact prior features and queries at the instance level and click, respectively. The HPQuery of addition, replacement, and concatenation operations can be expressed as:
@@
We will detail the effectiveness of these feature fusion methods in the experimental section.
Figure 3: Structure of UVE. Hybrid prior embedding enhances vector representation. Intra-instance and inter-instance attention mechanisms are employed to refine local features and capture global context effectively.
3.2 Architecture of UVE
Inspired by BERT [4] on text-related tasks, we propose a unified vector encoder (UVE) by analogizing vector points to words and vector map elements to sentences (see Figure 3). And we provide a pre-training strategy to improve the encoding and noise reduction ability of UVE.
Hybrid prior embedding.
As shown in Figure 3, for the given prior map $M_{prior}$, we first construct the input for UVE, termed as Hybrid Prior Embedding (HPE). We use the method from [30] for obtaining point position embedding and direction embedding for both $(x,y)$ and $(v_{x},v_{y})$. We concatenate these two embeddings as the point-level embedding. To obtain the feature representation of the entire vector, we add a special [VEC] token at the beginning of each vector and use the embedding of this token as the instance-level embedding. Additionally, we introduce learnable instance embedding and learnable type embedding to distinguish vector instances. To ensure the order of points, we also introduce 2D learnable position embedding. Finally, these embeddings are aggregated to form the HPE, which then serves as the input embedding for UVE.
Dual encoding mechanism.
Due to the limited information content of vector points themselves, learning the concept of vector instances is very challenging for the model, requiring more interactions among the points within the vectors to enhance each point’s perceptual abilities towards others in the same vector. As a result, the UVE encoder utilizes $M$-layer intra-vector attention encoders and $N$-layer inter-vector attention encoders, using an attention masking mechanism to facilitate feature interactions within and between different vector instances.
3.3 Pre-training UVE: Position modeling
Due to the limited inference ability of online mapping model, there are errors in historical prediction maps. So we used position modeling to pre-train UVE to improve its encoding and noise reduction capabilities. specifically, there are two methods: random noise and masking (see Figure 4).
Noise & mask generator.
The noise is primarily categorized into segment-level and point-level. In point-level noise, random noise is added to $5\%$ of the vector points across the entire map. For segment-level noise, random noise is added to all vector points within a sub-vector segment, after randomly selecting $10\%$ of the map elements. Using $M_{org}$ to represent the original map elements (with the same data structure as $M_{prior}$), the process of adding Gaussian noise can be represented by the following formula:
$M_{org}^{*}=M_{org}\emph{[random\_index]}+\epsilon,$ (8)
Similar to adding noise to the points, the selection of points for adding the mask is also divided into point-level and segment-level. After selecting the points to add the mask, the coordinates of these points will be masked as $-1$.
[…]
Loss function.
We feed the vector map to UVE encoding after passing through the noise mask generator, then decode the coordinates of all vector points using MLP, and calculate mean euclide distance from real coordinates as supervision:
$\mathcal{L}=RMSE(P,~{}mlp(E_{uve}(M_{org}^{*}))),$ (10)
where $P=\{(x_{i},y_{i})\}_{i=0}^{k}$ represent all the point in $M_{org}$.
4 Experiments
4.1 Experimental settings
Method Modality Backbone Epoch Metric ped. div. bou. mean $\mathbf{\tau_{0.5m}}$
HDMapNet [15] C Effi-B0 30 IoU 18.7 40.6 39.5 32.9 -
# SD map Prior (Ours) C Effi-B0 30 IoU 22.2 42.0 41.5 35.2 -
# Online Local Map Prior (Ours) C Effi-B0 30 IoU 23.1 41.4 42.0 35.5 -
# HD map-EX* Prior (Ours) C Effi-B0 30 IoU 21.2 42.3 44.9 36.1 -
MapTR [19] C R50 24 AP 41.2 49.5 51.1 47.3 -
# P-MapNet [13] C R50 24 AP 43.7 50.9 53.5 49.4 -
StreamMapNet [37] C R50 24 AP 60.4 61.9 58.9 60.4 -
# SQD-MapNet [32] C R50 24 AP 63.0 65.5 63.3 63.9 -
MGMap [22] C R50 24 AP 61.8 65.0 67.5 64.8 -
MapTRv2 [20] C R50 24 AP 59.8 62.4 62.4 61.5 38.6
# SD map Prior (Ours) C R50 24 AP 61.7 65.5 66.4 64.5 42.7
# Online Local Map Prior (Ours) C R50 24 AP 63.6 67.3 66.2 65.7 44.7
# HD map-EX* Prior (Ours) C R50 24 AP 61.5 65.6 68.3 65.1 44.3
Table 1: Performance of various online mapping models with different prior maps on nuScenes within a $60m\times 30m$ perception range. “C” denotes camera input. The SD map is sourced from OpenStreetMap [10]. The online local map refers to the historical prediction results. As the nuScenes dataset lacks existing HD maps, we followed the approach of MapEX [29] to create an HD map-EX*, simulating existing HD maps by removing pedestrian crossings and lane dividers, while retaining only the road boundaries.
@@ 4.1 Experimental settings
We mainly validate our method on the popular nuScenes [1] and Argoverse 2 [33] following the standard setting of previous methods [25, 19]. NuScenes contains a total of 1,000 scenes, each with an approximate duration of 20 seconds. The data was collected utilizing a 32-beam LiDAR operating and six cameras offering a 360-degree field of view. Argoverse 2 includes 1000 logs, each offering 15 seconds of 20Hz RGB images captured from 7 cameras, along with a 3D vector map.
Furthermore, to demonstrate that PriorDrive is plug-and-play and widely applicable to other vector tasks, we conducted a topology inference task on OpenLane-V2 [31] dataset. Subset A, developed by Argogorse 2 [33], includes seven views as input, along with an additional SD map.
@@
For the vectorized results, we follow the standard metric used in previous works [15, 25, 19]. The perception range along the direction of vehicle travel are $[-15.0m,15.0m]$ for the $X$-axis and $[-30.0m,30.0m]$ for the $Y$-axis. And we adopt average precision (AP) evaluated the quality of map construction. We compute the $\mathrm{AP}_{\tau}$ under several Chamfer distance $D_{Chamfer}$ thresholds $(\tau\in T,T=\{0.5,1.0,1.5\})$, and then calculate the mean across all thresholds as the the final AP metric:

### v4

# PriorDrive: Enhancing Online HD Mapping with Unified Vector Priors

@@ Abstract
High-Definition Maps (HD maps) are essential for the precise navigation and decision-making of autonomous vehicles, yet their creation and upkeep present significant cost and timeliness challenges. The online construction of HD maps using on-board sensors has emerged as a promising solution; however, these methods can be impeded by incomplete data due to occlusions and inclement weather, while their performance in distant regions remains unsatisfying. This paper proposes PriorDrive to address these limitations by directly harnessing the power of various vectorized prior maps, significantly enhancing the robustness and accuracy of online HD map construction. Our approach integrates a variety of prior maps uniformly, such as OpenStreetMap’s Standard Definition Maps (SD maps), outdated HD maps from vendors, and locally constructed maps from historical vehicle data. To effectively integrate such prior information into online mapping models, we introduce a Hybrid Prior Representation (HPQuery) that standardizes the representation of diverse map elements. We further propose a Unified Vector Encoder (UVE), which employs fused prior embedding and a dual encoding mechanism to encode vector data. To improve the UVE’s generalizability and performance, we propose a segment-level and point-level pre-training strategy that enables the UVE to learn the prior distribution of vector data. Through extensive testing on the nuScenes, Argoverse 2 and OpenLane-V2, we demonstrate that PriorDrive is highly compatible with various online mapping models and substantially improves map prediction capabilities. The integration of prior maps through PriorDrive offers a robust solution to the challenges of single-perception data, paving the way for more reliable autonomous vehicle navigation. Code is available at https://github.com/MIV-XJTU/PriorDrive
@@
Figure 1: Overview of PriorDrive. PriorDrive seamlessly integrates diverse vectorized prior maps into existing online mapping frameworks, making final predictions more complete and accurate than those generated without priors. The optimized predictions can be uploaded to cloud servers for other vehicles to download and use as prior maps.
@@
footnotetext: Work done during the internship at Amap, Alibaba Group.footnotetext: Equal contribution.footnotetext: Corresponding author. This work was supported by the National Natural Science Foundation of China No. 62572385 and the Fundamental Research Funds for the Central Universities No. xxj032023020.
Introduction
@@
High-Definition Maps (HD maps) are essential for the accurate navigation and decision-making of autonomous vehicles, providing detailed vectorized representations of road elements (HDmaps-survey; Gao2020VectorNetEH; yao2024waterscenes; yao2025exploring; lu2025differentiable). Despite their significance, the traditional methods for creating and maintaining HD maps are often costly and labor-intensive. These methods can result in outdated maps that struggle to keep pace with the rapidly changing urban environments. In response, there is increasing interest in online HD map construction (BEMAPNET; PivotNet; MAPVR; yao2025usvtrack), where maps are generated in real-time using on-board sensors. Although this approach reduces costs, it also introduces new challenges, particularly concerning incomplete and error-prone data caused by environmental occlusions (see red circle in Figure 1) or inclement weather conditions, coupled with unsatisfactory performance in distant regions.
Using prior maps to alleviate the problem of online mapping is a promising solution. Specifically, there are three common types of prior maps: 1) Standard Definition Maps (SD maps), 2) existing offline outdated HD maps, and 3) online historical prediction local maps, as illustrated in the upper left corner of Figure 1. SD maps provide crucial long-distance centerline skeletons, but with uncertain accuracy compared to HD maps. While existing HD maps (HD map-EX) offer high accuracy, their infrequent updates may lead to outdated information that fails to fully reflect current road conditions. Historical predicted local maps can offer insights by incorporating previous observations, but being a single mapping result, they may not guarantee the completeness and accuracy of the prior map.
Generally, there are two ways to encode vectors: 1) rasterized, 2) vectorized. Recent studies (Xiong2023NeuralMP; Jiang2024PMapNetFM; li2024towards1; li2025hyfacialhybrid) have attempted to use rasterized or vectorized prior maps to alleviate the problem of online mapping. However, existing prior-based methods (Xiong2023NeuralMP; Jiang2024PMapNetFM; MAPEX; 10.1117/12.3049486; li2025lion) face several limitations. For instance, P-MapNet (Jiang2024PMapNetFM), NMP (Xiong2023NeuralMP) and HRMapNet (zhang2024hrmapnet) leverage rasterized SD maps or historical prediction maps. The rasterized representation, limited by resolution, is lossy and redundant while lacking the detail necessary to capture vectorized instance-level information and express the type and direction of map elements effectively. Furthermore, P-MapNet and HRMapNet require complex post-processing to convert vectorized maps into rasterized maps, which also do not align with the native vector storage format of maps. On the other hand, MapEX (MAPEX) exclusively relies on vectorized outdated HD maps, which struggle to reflect real-time changes in road structures, due to low update frequencies, making them unreliable for current navigation needs. Moreover, the encoding scheme of MapEX is overly simplistic, thereby failing to effectively capture instance-level information within vector representations.
It is worth mentioning that encoding vectors is not a straightforward task. There are several main challenges: 1) Different types of vector maps may contain varying vector types (e.g., points, lines, planes), 2) Vector lengths can be fixed or variable (see blue circle in Figure 1), and 3) Alignment and inconsistency issues exist between different vector maps (see purple circle in Figure 1). Previous methods were limited to efficiently encoding only a single type of prior map, but the complementary information contained across different prior map categories. For instance, NavMap (schmidt2023navmap) omitted road segments with variable lengths when integrating HD maps and SD maps, resulting in the loss of valuable information. Furthermore, previous methods generally encode only vector positions and categorical attributes, lacking the capability to capture finer-grained information such as direction, geometric shape, and topology.
In this paper, we overcome these challenges and introduce PriorDrive, a novel framework designed to directly and unifiedly integrate diverse vectorized prior maps into various mapping models, fully leveraging their respective strengths and complementary information. We propose a Unified Vector Encoder (UVE) that encodes various vector data and captures fine-grained vector information. It can extract fixed-length instance-level and point-level features from variable-length vectors. Additionally, we introduce for the first time a point-level and segment-level pre-training paradigm specifically tailored for vector data, aiming to enhance the model’s understanding and representation of vector. Concurrently, we propose a hybrid prior representation (HPQuery) to represent all elements, facilitating the integration of unified vector features into mapping models. Experiments on nuScenes, Argoverse 2 and OpenLane-V2 demonstrate that PriorDrive is plug-and-play, seamlessly applicable to diverse online mapping models and vector tasks with significant performance improvements.
@@
  • We introduce a unified vector encoder that effectively captures diverse vector attributes and enables encoding of variable-length vectors into fixed-length features through fused prior embedding and dual encoding mechanism.
@@
  • We first proposed a point-level and segment-level pre-training paradigm for vector data, effectively capturing prior distribution and mitigating noise in historical maps.
@@
  • We propose a hybrid prior representation (HPQuery) to represent all elements and a PriorDrive framework with various vector prior maps to address the limitations of single-perception. Our comprehensive evaluation demonstrates that PriorDrive is plug-and-play and significantly enhances the online mapping models.
Related Work
Online Vectorized HD Map Construction
@@
Online vectorized HD map construction was initially treated as a segmentation task (bevformer; zhang2024self; sun2023large; guo2025depth; cui2025efficient). To construct a vectorized HD map, HDMapNet (Li2021HDMapNetAO) requires complex post-processing of pixel-level rasterized maps. VectorMapNet (Liu2022VectorMapNetEV) introduced a coarse-to-fine, two-stage network leveraging keypoint representations. MapTR (liao2022maptr) advanced this concept by modeling point sets in a permutation-equivalent manner using a DETR-like (detr; 9681188; zeng2025janusvln; xiao2025worldenv; jin2025reasoning) one-stage network. Subsequent studies (HIMAP; sun2025yolov4svm; zeng2024driving; lan2025contextual; lan2025mappo) have introduced various insightful approaches for further improvement. However, these methods rely solely on single-source perception data from vehicle-mounted sensors, limiting their effectiveness in challenging scenarios such as occlusion, inclement weather conditions or distant regions.
@@
The release of OpenLane-V2 (wang2023openlanev2) dataset has sparked growing interest in topology reasoning within driving scenes. This task focuses on recognizing the topologies among lanes and lanes with traffic elements. TopoNet (li2023toponet) leverages GNN to update lane representation and lane topology. TopoMLP (wu2023topomlp) utilizes PETR (Liu2022PETRPE) for centerline detection and employs simple MLPs to predict the relationships. In this paper, we also apply PriorDrive to topological inference tasks, demonstrating that PriorDrive is plug and play and widely applicable to other vector tasks.
Online Mapping Based on Prior Maps
@@
Recent research has explored the integration of prior maps to enhance the performance of online mapping models. P-MapNet (Jiang2024PMapNetFM) encodes Standard Definition Maps (SD map) as an additional conditional branch and employs a masked autoencoder to capture the prior distribution of HD maps. NMP (Xiong2023NeuralMP) introduces a global neural map prior that self-updates, thereby improving the performance of local map inference. MapEX (MAPEX) utilizes existing HD maps and refines the query-based map estimation model’s matching algorithm. Although these advancements have achieved considerable results, many studies have overlooked the potential of online historical prediction maps as a source of prior information. Moreover, previous methods (zhang2024hrmapnet; Ma2024RoadPainterPA; wan2025driving; yuan2025unimapgen; li2025multi) typically could only utilize a single type of prior map. Our work can uniformly encode various types of vector prior maps to enhance current perception data, fully leveraging the complementary advantages among different prior maps.
Pretrained Methods Based on Mask Modeling
@@
In the fields of NLP and CV, masked pre-training has proven to be an effective strategy for self-supervised representation learning. In NLP, models like BERT (betr; Liu2019RoBERTaAR; Lan2019ALBERTAL; zeng2025FSDrive; li2023ultrare) use masked language modeling to predict randomly masked tokens within a bidirectional text context. Similarly, methods such as MAE (MAE; Liu2022MixMAEMA; liang2025persistent; dai2025unbiased; yin2025knowledge) mask random patches of input images and reconstruct them based on the remaining unmasked patches. Diverging from these approaches (zhang2025molebridge; ZHANG2025103888; wei2025copeft; wei2026infocom; li2025cogvla), which are predominantly tailored for text and image data, we introduce a pre-training paradigm specifically designed for vector data. Our method aims to learn the prior distribution of vector representations, offering a new avenue for pre-training in the context of vectorized data.
Proposed Method: PriorDrive
Problem Formulation
@@
The goal of online mapping is to construct a local HD map using on-board sensor observations, such as the surrounding camera images. The online mapping process typically involves two main modules: the Bird’s-Eye View (BEV) encoder and the map decoder. The BEV encoder generally includes a Perspective View (PV) feature extractor and a module for transforming PV features into BEV features. The map decoder usually comprises a Fully Convolutional Network (Shelhamer2014FullyCN; hu2024tmff; 202508.0462; huang2024ar) or a query-based module similar to DETR (detr; xie2025seqgrowgraph; tang2025dissecting; tang2025ocrt), which outputs either rasterized perception results or vectorized map elements.
Figure 2: Structure of UVE. Fused prior embedding enhances vector representation. Intra-instance and inter-instance attention refine local features and capture global context effectively.
@@
where $\mathcal{P}$ represents the predicted map elements, $Q$ denotes the learnable query, and optopt indicates that this learnable query is optional, which is required in query-based models but not in other models.
Architecture of UVE
Inspired by BERT (betr; he2025social; rs17132267; ren2025digital) on text-related tasks, we propose a unified vector encoder (UVE) by analogizing vector points to words and vector elements to sentences (see Figure 2).
@@
Our proposed UVE serves as a unified vector encoder that can directly encode various vector data information, such as the position ($x$, $y$), direction ($v_{x}$, $v_{y}$), and type of points ($c$). Using $p=\left[x,y,v_{x},v_{y},c\right]$ to represent the point, a vector $v$ typically consists of an ordered set of a variable number of points, $v=\left[p_{1},p_{2},\ldots,p_{n}\right]$, where $n$ varies. The map prior $M$ consists of a set of vectors $M=\{v_{1},v_{2},\ldots,v_{m}\}$, with $m$ being variable. When employing different prior maps, all of prior maps were integrated into a cohesive whole, $M_{prior}=\{M_{1},M_{2},\ldots,M_{t}\}$. Using $E_{uve}$ to represent the UVE model, the extraction process of the features of the prior map, denoted as $f_{prior}$, can be represented as:
@@
Fused Prior Embedding.
As shown in Figure 2, for the given prior maps $M_{prior}$, we first construct the input for UVE, termed as Fused Prior Embedding (FPE). We use the method from (TancikSMFRSRBN20; zeng2025lens; lin2025plan; yang2025unleashing) for obtaining point position embedding and direction embedding for both $(x,y)$ and $(v_{x},v_{y})$. We concatenate these two embeddings as the point-level embedding. To obtain the fixed-length feature of the entire vector, we add a special [VEC] token at the beginning of each vector and use the embedding of this token as the instance-level embedding. Additionally, we introduce learnable instance embedding and learnable type embedding to distinguish vector instances. To ensure the order of points, we also introduce 2D learnable position embedding. Finally, these embeddings are aggregated to form FPE.
Dual Encoding Mechanism.
Due to the limited information content of vector points themselves, learning the concept of vector instances is very challenging, requiring more interactions among the points within the vectors to enhance each point’s perceptual abilities towards others in the same vector. Therefore, UVE encoder utilizes $M$-layer intra-vector attention and $N$-layer inter-vector attention, using an attention masking mechanism to facilitate feature interactions within and between different vector instances. Each mask in the intra-vector encoder has a value of 1 at its own position and 0 elsewhere, while the inter-vector encoder employs a mask matrix filled entirely with 1s.
When using different prior maps and multiple online local maps, there are alignment problems between different prior maps and potential inconsistencies between online local maps generated across different trips or different timestamps in one trip. Our UVE mitigated this challenge by employing fused prior embedding and dual encoding mechanism to enhance the model’s ability to learn instances, supporting adaptive focus on the superior map elements in different prior maps. The experiments presented in Table 4 substantiate this capability.
Pre-training UVE: Position Modeling
Due to the limited inference ability of online mapping models, there are errors in historical prediction maps. So we used position modeling to pre-train UVE to improve its encoding and noise reduction capabilities. Specifically, there are two methods: random noise and masking (see Figure 3).
Noise & Mask Generator.
The noise is primarily categorized into segment-level and point-level. In point-level noise, random noise is added to $5\%$ of the vector points across the entire map. For segment-level noise, random noise is added to all vector points within a sub-vector segment, after randomly selecting $10\%$ of the map elements. Using $M_{org}$ to represent the original map elements (with the same data structure as $M$), the process of adding Gaussian noise can be represented by the following formula:
$M_{org}^{*}=M_{org}\emph{[random\_index]}+\epsilon,$ (4)
Similar to adding noise, the selection of points for adding the mask is also divided into point-level and segment-level. After selecting the points to add the mask, the coordinates of these points will be masked as 1-1.
$M_{org}\emph{[random\_index]}=Mask.$ (5)
Loss Function.
We feed vector map to UVE encoding after passing through the noise mask generator, then decode the coordinates of all points using MLP, and calculate mean euclide distance from GT coordinates as supervision:
$\mathcal{L}=RMSE(P,~mlp(E_{uve}(M_{org}^{*}))),$ (6)
where $P=\{(x_{i},y_{i})\}_{i=0}^{k}$ represent all the points in $M_{org}$, $RMSE$ on behalf of the root mean square error.
Figure 4: Our Integration Method. For query-based models, we enable interaction between the prior features and queries at both instance-level and point-level, ensuring a more nuanced integration. For the other models, we integrate prior features directly with the BEV features.
Integrated Approaches to BEV and HPQuery
We respectively proposed different integration methods for the query-based/non-query-based approaches to integrate the prior map features into the online map construction model. For methods without a learnable query $Q$, such as HDMapNet (Li2021HDMapNetAO), we directly reshape the prior feature $f_{prior}$ and use $Deconv$ upsampling to match the shape of the BEV feature $f_{bev}$. Then connect it with the BEV feature $f_{bev}$ and aligned by $Conv$, as shown in Figure 4 left.
@@
For methods based on learnable query $Q$, such as the vectorized model MapTR series (maptrv2), we propose HPQuery of three operations: addition, replacement, and concatenation, as shown in Figure 4 right. Specifically, $Q$ is typically composed of two parts, $Q=\{q_{ij}\}_{i=0,j=0}^{m,n}=\{q_{i}^{ins}+q_{j}^{pt}\}_{i=0,j=0}^{m,n}$, $q_{i}^{ins}$ and $q_{i}^{pt}$ representing the learnable queries at the instance-level and point-level respectively. The features $f_{prior}$ of prior maps are also composed of instance-level features and point-level features, $f_{prior}=\{(f_{i}^{ins},f_{j}^{pt})\}_{i=0,j=0}^{m^{\prime},n^{\prime}}$. So we interact prior features and queries at the instance-level and point-level, respectively. The HPQuery of addition, replacement, and concatenation operations can be expressed as:
@@
We will detail the effectiveness of these feature fusion methods in supplementary material.
Experiments
Experimental Settings
@@
We mainly validate our method on the popular nuScenes (Caesar2019nuScenesAM) and Argoverse 2 (Argoverse2) following the previous methods (zhang2024hrmapnet; maptrv2; zhou2025hademif; zhou2025valuing). NuScenes contains a total of 1,000 scenes, which was collected utilizing a 32-beam LiDAR operating and six cameras offering a 360-degree field of view. Argoverse 2 includes 1000 logs, which was captured from 7 cameras, along with a 3D vector map.
To demonstrate PriorDrive is plug-and-play and widely applicable to other vector tasks, we conducted topology inference on OpenLane-V2 (wang2023openlanev2) Subset A.
@@
For vectorized results, we follow the standard metric used in previous works (Liu2022VectorMapNetEV; maptrv2; li2025revolutionizing; li2025knowledge). The perception range along the direction of vehicle travel is [15.0m,15.0m][-15.0m,15.0m] for $X$-axis and [30.0m,30.0m][-30.0m,30.0m] for $Y$-axis. We compute the average precision (AP) $\mathrm{AP}_{\tau}$ under several Chamfer distance $D_{Chamfer}$ thresholds $(\tau\in T,T=\{0.5,1.0,1.5\})$, and then calculate the mean across all thresholds as the final AP metric: